Suppose you have 0.664 mol of an ideal gas in a 0.454 L container at 39.97 atm.
a) Sketch a pressure versus volume graph for this ideal gas, either using the graph paper on the next page or in a spreadsheet such as Excel. Have the graph go out to 4 L and 70 atm. Use at least 6 points to draw this curve.
b) On the graph from part b, use a dotted line to sketch the approximate curve that you might expect to see for a real gas with strong repulsive intermolecular interactions at high pressures (i.e. above 10 atm).
Answer #1

As we know that

PV = nRT

Thus; T = PV/nR = (39.91 atm x 0.454 L) / (0.664 mol x0.08205 ((L x atm) / (mol x K )) = 333 K

So, as per the formula PV = nRT

P = nrtT/V = (0.664 mol x 0.08205 ((L x atm) / (mol x K )) x 333K ) / V = 18.1422396‬ / V

Thus the graph would be: